OpenStudy (anonymous):

guys pls solve.. hehe 1+ cos2x / sin2x

OpenStudy (lgbasallote):

is that supposed to read \[\frac{1 + \cos (2x)}{\sin (2x)}\] or no?

hartnn (hartnn):

\[\huge{\frac{1+\cos 2x}{\sin 2x}}\] right?

OpenStudy (anonymous):

yes

hartnn (hartnn):

\[\cos 2x=2\cos^2 x -1\] \[\sin 2x= 2 \sin x \cos x\] does this help?

OpenStudy (anonymous):

the answer is cotx but i dont know how did it come up with that

hartnn (hartnn):

yes it is,just put values of cos 2x and sin 2x i gave in that fraction,and cancel out terms,u will get cot x...

OpenStudy (anonymous):

ahh yey! thank you! :D

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

wait i got confused ,, hehehe.. can u xplain it again? if thats alright with you.. :))

hartnn (hartnn):

sure,which part?

hartnn (hartnn):

\[\frac{1+\cos 2x}{\sin 2x}=\frac{2\cos^2 x}{2\sin x \cos x}=\frac{2cosx \cos x}{2\sin x \sin x}=\frac{\cos x}{\sin x}=cotx\]

OpenStudy (lgbasallote):

i think you made a typo there...

hartnn (hartnn):

yup ,sorry. \[\frac{1+\cos 2x}{\sin 2x}=\frac{2\cos^2 x}{2\sin x \cos x}=\frac{2cosx \cos x}{2\sin x \cos x}=\frac{\cos x}{\sin x}=cotx\]

OpenStudy (anonymous):

\[2 \cos ^{2}x\]

OpenStudy (anonymous):

how did you get that?

hartnn (hartnn):

the formula gor cos2x is 2cos^2 x -1 so 1+cos2x= 2 cos^2 x

OpenStudy (anonymous):

i thought it should b \[2\cos ^{2}x - 1\]

OpenStudy (anonymous):

ahh ok.. iknow now tnx alot

hartnn (hartnn):

thats the formula for cos 2x,but u have +1 in the numerator also

OpenStudy (anonymous):

how about this.. wait

OpenStudy (anonymous):

\[\frac{ \cos ^{2} \theta - \sin ^{2}\theta}{\sin \theta \cos \theta }\]

hartnn (hartnn):

another representation of cos 2x is cos^2 x-sin^2 x so numerator directly is cos 2x for denominator,u can uswe,sin 2x =2 sin x cos x

hartnn (hartnn):

ok?

hartnn (hartnn):

the answer would be then 2 cot 2x

OpenStudy (anonymous):

tnx